OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20

OpenStudy (anonymous):

solve algebraically...

OpenStudy (anonymous):

|2x + 6|=20+4=24 2x+6=24 2x=24-6=18 x=18/2=9 x=9

OpenStudy (anonymous):

check for the value of x... particularly the absolute value of 2x+6...

OpenStudy (anonymous):

|2(9)+6|=|18+6|=|24|=24... therefore x=9 is correct, actually the absolute of any number is always the positive number...

OpenStudy (anonymous):

hope you understand it...